OpenStudy (dls):

\[\LARGE \lim_{x \rightarrow \infty} (\frac{x^2+5x+3}{x^2+x+3})^x\]

OpenStudy (anonymous):

is it ^x or * x

OpenStudy (dls):

^x

OpenStudy (anonymous):

Only for numerator?

OpenStudy (dls):

no..whole

OpenStudy (anonymous):

Can we apply L-Hospital's Rule..

OpenStudy (anonymous):

@DLS

OpenStudy (dls):

yes

OpenStudy (anonymous):

Then...Differentiate Numerator and denominator Seperately..)

OpenStudy (dls):

e^4 e^2 e^3 1

OpenStudy (dls):

@Yahoo! didnt get the ans

OpenStudy (experimentx):

Use this \[ \huge a = e^{\log a}\] then use L'Hopital

OpenStudy (dls):

:O what in the world is that??

geerky42 (geerky42):

Just use your logic. We know that \(\large lim_{x \rightarrow \infty}\dfrac{x^2}{x^2} = 1\) So \(\large \lim_{x \rightarrow \infty}\left(\dfrac{x^2+5x+3}{x^2+x+3}\right)^x = \lim_{x \rightarrow \infty}1^x = \boxed{1}\)

OpenStudy (dls):

logic fails answer isnt 1

geerky42 (geerky42):

And how do you know?

OpenStudy (anonymous):

@geerky42 ^x is only for numerator

OpenStudy (dls):

its for whole of it

OpenStudy (dls):

not only num

OpenStudy (dls):

@geerky42 relax..answer isn't 1..

geerky42 (geerky42):

Well, how do you know?

OpenStudy (anonymous):

Lol....u said "no..whole" hm..hm

OpenStudy (dls):

lol i said no..not num..whole

OpenStudy (anonymous):

oh...then...i was mistaken..)

OpenStudy (experimentx):

@Yahoo! try that

geerky42 (geerky42):

And what in tarnation kind of class are you taking, anyway?

OpenStudy (anonymous):

but my answer too ended up in 1..@experimentX

geerky42 (geerky42):

How do you know answer isn't one? @DLS

OpenStudy (dls):

i have the solution? :/

geerky42 (geerky42):

Why do you need our help if you have solution?

OpenStudy (dls):

I DIDNT UNDERSTAND IT ?

OpenStudy (dls):

correct

OpenStudy (dls):

@geerky42 hope that satisfied you.

geerky42 (geerky42):

You mean answer key, not solution...

geerky42 (geerky42):

Do you have to show your work?

OpenStudy (dls):

they have done weird things \[\LARGE \lim_{x \rightarrow \infty} [(1+\frac{4x+1}{x^2+x+2})^{\frac{x^2+x+2}{4x+1}}]^\frac{(4x+1)x}{x^2+x+2}\]

geerky42 (geerky42):

What in tarnation kind of class are you taking?

geerky42 (geerky42):

Since it appears that we cannot solve it, just use WolframAlpha. Good luck...

OpenStudy (dls):

Don't use "We".

OpenStudy (experimentx):

this is straight forward L'Hopital rule put that base in log an raise it to e .. then you get type inf x 0 ... use L'hopital rule couple of times, you should get answer

OpenStudy (dls):

still ends up complicated :/

OpenStudy (experimentx):

how to do you evaluate \[ \lim_{x \to \infty} x \log\left( x^ 2 +5x+3 \over x^2 + x+ 3\right)\]

OpenStudy (dls):

wait i guess i got it :O

OpenStudy (dls):

thanks! :D

OpenStudy (experimentx):

for all functions of type \[ \huge f(x)^{g(x)} = e^{\log f(x)^{g(x)}} = e^{g(x) \log f(x)}\]